OpenStudy (anonymous):

The length of a rect. is 6 more than 4 times the width. the perimete is 62. find the dimensions of the rectangle.

OpenStudy (anonymous):

i keep getting it wrong

OpenStudy (anonymous):

x should equal 5 but i got 29

OpenStudy (apoorvk):

wait. i made a mistake. didn't read it carefully.

OpenStudy (anonymous):

yeah it should be 4x+6

OpenStudy (anonymous):

or 6(4x)

myininaya (myininaya):

more means add

OpenStudy (apoorvk):

let the rectangle's width be 'x'. Then, length = 6 more than 4 times the width= 4x +6 Perimeter = 62 = 2(length + breadth) put in the values of length and the breadth on terms of 'x', and find 'x' by solving the equation.

OpenStudy (anonymous):

yeah so 4x+6 but i tried that and it didnt wrk

myininaya (myininaya):

so L is 6 more than 4 times the width is L = 6 + 4 W

myininaya (myininaya):

Then use the formula for perimeter since the perimeter is given 2L+2W=62

OpenStudy (anonymous):

i got 11.2 and thats not the answer ok ll try that

OpenStudy (anonymous):

IT WORKED!! thanks muninaya!

OpenStudy (apoorvk):

actually you will get x=5 alright, if you do this correctly.